Solve the system of equations using the addition (elimination) method.
If the answer is a unique solution, present it as an ordered pair: (x, y). If not, specify whether the answer is “no solution” or “infinitely many solutions.”
4x – 3y = 1
-12x + 9y = 5

Answer by padmameesala(42)

4x – 3y = 1
-12x + 9y = 5 that is 4x-3y = -5/3 by dividing the equation by -3
so both the equations have same slope and different y-intercepts
that is the two lines are paralle; and they dont have point of intersection
so the given system of equations have no solution
hope you will understand it